Grade 12th passIIT JEE Entrance Exam

Consider a nonconducting ring of radius r and mass m which has a charge q distributed uniformly over it. the disc is rotated about its axis with an angular speed w (w=2*22/7).find magnetic moment of the disc

shivam
9 Years agoGrade 12th pass

1 Answer

Arun
6 Years ago
(a) i = q/t = q/(2π / ω) = qω/2π (b) μ = n ia = i A [∴ n = 1] = qωπ r^2/2π = qωr^2/2 (c) μ = qωr^2/2, L = Iω = mr^2 ω, μ/L = qωr^2/2mr^2ω = q/2m ⇒ μ = (q/2m)